Help me please, I do not know how to solve these types of equations,
0.5x+0.4y=-1
0.3x+0.6y=-0.6

Answer by checkley71(8403)

YOU NEED TO ELIMINATE ONE OF THE UNKNOWN TERMS BY ADDITION OR SUBTRACTION. THUS:
MULTIPLY THE FIRST EQUATION BY -3 & THE SECOND EQUATION BY 5 & ADD.
-1.5-1.2Y=3
1.5X+3Y=-3
------------------------
1.8Y=0
Y=0 ANSWER. NOW SUBSTITUTE 0 FOR Y & SOLVE FOR X
.5X+0*Y=-1
.5X=-1
X=-1/.5
X=-2 ANSWER.
PROOF
.3*-2+.6*0=-.6
-.6=-.6